# FEDONet : Fourier-Embedded DeepONet for Spectrally Accurate Operator Learning

### v1

@@ Abstract
Deep Operator Networks (DeepONets) have recently emerged as powerful data-driven frameworks for learning nonlinear operators, particularly suited for approximating solutions to partial differential equations (PDEs). Despite their promising capabilities, the standard implementation of DeepONets, which typically employs fully connected linear layers in the trunk network, can encounter limitations in capturing complex spatial structures inherent to various PDEs. To address this, we introduce Fourier-embedded trunk networks within the DeepONet architecture, leveraging random Fourier feature mappings to enrich spatial representation capabilities. Our proposed Fourier-embedded DeepONet, FEDONet demonstrates superior performance compared to the traditional DeepONet across a comprehensive suite of PDE-driven datasets, including the two-dimensional Poisson equation, Burgers’ equation, the Lorenz-63 chaotic system, Eikonal equation, Allen–Cahn equation, Kuramoto–Sivashinsky equation, and the Lorenz-96 system. Empirical evaluations of FEDONet consistently show significant improvements in solution reconstruction accuracy, with average relative $\ell_{2}$ performance gains ranging between 2-3x compared to the DeepONet baseline. This study highlights the effectiveness of Fourier embeddings in enhancing neural operator learning, offering a robust and broadly applicable methodology for PDE surrogate modeling.
@@ 1 Introduction
Partial differential equations (PDEs) underpin models ranging from turbulent-flow dynamics and heat diffusion to electromagnetics and biological transport. Classical numerical solvers—finite-difference, finite-element and spectral methods—provide high-fidelity solutions but incur a prohibitive computational burden when faced with high-dimensional, multiscale or stiff systems [1, 2, 3]. The quest for rapid surrogates has a long pedigree: universal-approximation theorems for neural networks [4, 5, 6, 7], reduced-order techniques based on proper-orthogonal decomposition (POD) [8, 9], kernel–Gaussian-process surrogates [10], and mesh-free radial-basis schemes [11, 12]. Yet, translating these ideas into *function-to-function* (operator) mappings that respect arbitrary discretisations remained elusive.
The modern wave of operator learning addresses this gap by training neural networks that act directly on infinite-dimensional function spaces and generalise across grid resolutions. Deep Operator Networks (DeepONets) [13] and the broader Neural Operator family [14, 15, 16] establish a branch–trunk architecture with rigorous operator–universal-approximation guarantees. Fourier Neural Operators (FNOs) extend this paradigm through global spectral convolutions, delivering impressive speed-ups on parametric PDEs [17]. Subsequent variants exploit multiwavelets [18, 19], convolutional kernels [20], multi-grid tensorisation [21, 22] and geometry-aware deformations [23] to boost resolution or handle complex domains.
Beyond kernel choices, researchers have tackled discontinuities [24], derivative supervision [25], irregular meshes via graph trunks [26], active multi-resolution sampling [27], and U-Net/wavelet hybrids [28, 29]. Random Fourier features have proved especially effective at encoding high-frequency structure in standard deep nets [30], yet remain under-explored inside operator networks whose trunk is typically a plain fully connected multilayer perceptron (MLP).
To incorporate governing laws and uncertainty, Physics-Informed Neural Networks (PINNs) [31, 32] evolved into Physics-Informed Neural Operators (PINO) [33, 34] and their variational or pseudo-physics relatives [35, 36]. Complementary efforts embed operators in latent or invertible spaces for data efficiency and bidirectional inference [37, 38, 39, 40].
Transformer-style attention has recently been adopted for operator learning [41, 42, 43, 44, 45] and for geometry generalisation [46]. Such architectures underpin landmark climate-scale surrogates such as FourCastNet [47] and its amortised FNO successor [48].
Despite the rich taxonomy above, the trunk of conventional DeepONets, charged with encoding spatial/temporal coordinates, remains a shallow MLP. This design struggles to capture the oscillatory or multi-scale patterns ubiquitous in nonlinear PDE solutions, leading to sub-optimal sample efficiency and accuracy. Recent trunk alternatives (e.g. graph [26] or wavelet U-Net [28]) introduce architectural complexity or domain-specific preprocessing.
We propose a Fourier-Embedded DeepONet (FEDONet) that replaces the raw coordinates fed into the trunk with a compact random Fourier-feature map [30]. The embedding supplies an implicit spectrum of sinusoidal basis functions, endowing the trunk with a richer hypothesis space at negligible parameter cost. FEDONet preserves the original branch–trunk separation, is plug-and-play with existing DeepONet codebases, and adds no runtime overhead beyond a single matrix multiplication.
  1. (i)
    We formulate the Fourier embedding of trunk coordinates, derive its effect on the operator neural tangent kernel, and show that it strictly enlarges the approximation class of DeepONets.
  2. (ii)
    We deliver a thorough benchmark across seven PDE families—steady Poisson, viscous Burgers, Allen–Cahn, Kuramoto–Sivashinsky, Eikonal, Lorenz–63 and Lorenz–96—demonstrating 48.93% average reduction in relative $\ell^{2}$ error versus a tuned vanilla DeepONet.
  3. (iii)
    We analyse hyperparameter sensitivity (mapping size and scale) and empirically confirm that the Fourier trunk improves convergence speed without sacrificing discretisation invariance.
By injecting a principled spectral bias into the DeepONet trunk, FEDONet closes the performance gap to frequency-domain operators such as FNO while retaining the flexibility and locality advantages of the original architecture.
@@ 2.1 Fourier Embeddings as Spectral Preconditioners
Standard MLPs exhibit a well-known spectral bias — they learn low-frequency components faster than high-frequency ones [49]. This bottleneck is particularly harmful in operator learning problems involving sharp gradients, multiscale dynamics, or oscillatory fields.
To address this, we propose injecting fixed Fourier embeddings into the trunk network input. Given a coordinate $\zeta\in\mathbb{R}^{d}$, we define a randomized Fourier feature map:
@@
where $B\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ — enriching its spectral content.
The DeepONet prediction then becomes:
@@
effectively lifting $\zeta$ into a high-dimensional basis of sinusoidal functions. By Bochner’s theorem [50], such embeddings approximate shift-invariant kernels — implicitly kernelizing the trunk network. This expands its representational capacity, enabling better learning of high-frequency targets.
@@ 2.2 Training Objective
We adopt a supervised learning framework to train the proposed FEDONet model. The objective is to approximate a target operator $G$ by minimizing the prediction error over a dataset of paired input–output functions. Each dataset element consists of a function $u^{(i)}\in\mathcal{U}$ and corresponding evaluations of the output $s^{(i)}_{j}=G(u^{(i)})(\zeta^{(i)}_{j})$ at query locations $\zeta^{(i)}_{j}\in\mathcal{Y}$.
@@
Model parameters are optimized using mini-batch stochastic gradient descent with the Adam optimizer. Learning rate schedules are selected based on the problem characteristics to ensure stable and efficient convergence. Each mini-batch consists of randomly sampled function–query pairs $(u,\zeta)$, enabling scalable and efficient training across diverse domains.
@@ 3 Results
We created a synthetic dataset comprising numerical solutions of the two-dimensional (2D) Poisson equation, a fundamental PDE commonly employed in modeling steady-state processes such as heat conduction, electrostatics, and diffusion phenomena. The mathematical formulation of the Poisson equation is given by:
@@
accompanied by homogeneous Dirichlet boundary conditions:
$u(x,y)\big{|}_{\partial\Omega}=0.$ (10)
In this context, $u(x,y)$ denotes the solution field, while $f(x,y)$ represents a specified source term or coefficient.
To generate the dataset, we synthesized $N=10,000$ distinct instances of the coefficient function $f$ over a uniform spatial grid of resolution $128\times 128$. Each instance of the coefficient field was produced using Gaussian Random Fields (GRFs) characterized by specific spectral parameters, notably $\alpha=3$ and $\tau=3$, resulting in:
$f\sim\text{GRF}(\alpha=3,\tau=3).$ (11)
@@
Once these coefficient fields were created, we numerically solved the associated Poisson equation to obtain the corresponding solution fields $u(x,y)$. The numerical discretization employed a finite difference approach, utilizing a conventional five-point stencil approximation of the Laplacian operator. By discretizing the spatial domain $\Omega=[0,1]^{2}$ into a $128\times 128$ structured grid, the PDE was transformed into a linear algebraic system:
$A\mathbf{u}=b,$ (12)
where $\mathbf{u}\in\mathbb{R}^{16384}$ denotes the vectorized solution field and $b\in\mathbb{R}^{16384}$ is the vector representation of the source term $f$. The discretized Laplacian matrix $A$ is defined explicitly as:
[…]
with homogeneous Dirichlet boundary conditions enforced directly by setting boundary rows of $A$ as identity matrices and corresponding entries in $f$ to zero.
Ultimately, this methodology yielded 10,000 independent pairs of coefficient-source fields and their respective solutions, forming a robust dataset suitable for training, validating, and benchmarking neural operator models. The dataset facilitates a comprehensive evaluation of the trade-offs in computational efficiency, accuracy, and reconstruction performance of our proposed neural operator architectures.
Figure 2 presents a representative test sample from the two–dimensional Poisson dataset. The input forcing field is shown on the left, followed by the exact solution, the prediction of a vanilla DeepONet, and the corresponding point-wise error. The bottom row repeats the comparison for our Fourier-Embedded DeepONet (FEDONet).
  • Spatial reconstruction. FEDONet reproduces the lobed extrema, smooth gradients and corner behaviour of the ground truth almost indistinguishably, whereas the vanilla network exhibits visible blurring and slight spatial shifts of the high-magnitude regions. The residual of the vanilla model oscillates between $\pm 1.5\times 10^{-3}$, displaying high-frequency artefacts; FEDONet reduces this range to $\pm 3\times 10^{-4}$, a five-fold improvement that manifests as a low-amplitude, coherent pattern.
  • Energy spectra. Figure 3 plots the angle-integrated energy spectrum $E_{k}$ for four randomly chosen hold-out fields (samples 29, 15, 24 and 17). Both networks match the reference spectrum at large scales ($k\lesssim 5$), but qualitative differences arise in the inertial and dissipation ranges:
    1. (a)
      The vanilla DeepONet (purple) consistently over-estimates energy for $k\gtrsim 8$, producing a spectral “pile-up” that grows by nearly an order of magnitude in sample 24—mirroring the spurious small-scale oscillations seen in its error map.
    2. (b)
      FEDONet (red) tracks the ground truth (dark blue) almost perfectly up to the Nyquist limit, deviating by at most a factor of two in the far dissipation tail and thereby preserving the correct energy cascade.
Figure 2: Model reconstruction comparison for the 2D Poisson Dataset.
Together, the spatial and spectral diagnostics confirm that Fourier embeddings in the trunk endow DeepONets with the high-frequency resolution and stability required for accurate operator learning: FEDONet not only slashes point-wise errors but also eliminates the artificial energy plateau that plagues fully-connected trunks.
@@ 3.2 Burgers’ Equation
Figure 5: Best-performing test sample: Ground truth vs. DeepONet vs. FEDONet prediction.
@@
Figure 6: Temporal Evolution of Pointwise Error and Energy Spectrum: Best-Performing Sample vs. Ground Truth
@@
Figure 7: Training loss convergence for the 1D Burgers’ Equation: Comparison of Vanilla and Fourier Embedded DeepONets trained with varying batch sizes.
@@ 3.3 Lorenz–63 System
A dataset of 10,00010{,}000 trajectories is generated using the classic Runge–Kutta method (RK45), each evaluated at 1000 uniformly spaced time steps. The input to the DeepONet branch network is the scalar $x_{0}$, and the trunk network is queried at time coordinates $t\in[0,3]$. The goal is to accurately recover the full state trajectory, including lobe transitions and fine-scale oscillations characteristic of chaotic dynamics.
@@
Finally, Figure 12 compares the spectral content of the predicted and reference trajectories. The vanilla DeepONet fails to capture high-frequency modes, resulting in a smoothed trajectory that diverges from the correct dynamics. FEDONet, equipped with randomized Fourier features, preserves the broadband energy spectrum more accurately—an essential factor in recovering the fine-scale structure and temporal richness of chaotic systems.
@@ 3.4 Eikonal Equation
Figure 13: Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry. Relative $L^{2}$ error: $2.728\times 10^{-2}$ for DeepONet, $0.912\times 10^{-2}$ for FEDONet.
@@ 3.5 Lorenz-96 Chaotic System
To evaluate the performance of our DeepONet-based neural operator models, we utilize the Lorenz-96 (L96) system—a well-known chaotic dynamical system commonly employed for benchmarking data-driven models and surrogate learning techniques in climate modeling and nonlinear dynamics. The L96 system is defined by the set of coupled ordinary differential equations:
$\frac{dx_{i}}{dt}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,\quad i=1,2,\dots,N$ (27)
@@
with periodic boundary conditions: $x_{i-N}=x_{i}=x_{i+N}$. Here, $x_{i}$ denotes the state of the $i$-th variable, $N$ is the number of state variables, and $F$ is a constant forcing parameter that controls the nonlinearity and chaotic behavior of the system. For $F>8$, the system is known to exhibit strong chaos. In our experiments, we select $F=4.0$, which yields nonlinear but non-chaotic dynamics.
@@
The L96 system is integrated forward in time using the classical fourth-order Runge-Kutta (RK4) method. Given an initial condition $\mathbf{x}_{0}\in\mathbb{R}^{N}$, the RK4 method computes the solution at the next time step using:
$\mathbf{x}_{n+1}=\mathbf{x}_{n}+\frac{\Delta t}{6}(k_{1}+2k_{2}+2k_{3}+k_{4})$ (28)
where the intermediate slopes $k_{1},k_{2},k_{3},k_{4}$ are computed using the Lorenz-96 time derivative evaluated at successive RK stages, and $\Delta t=0.01$ is the time step.
To construct a comprehensive dataset for training and evaluating neural operator models, we generate an ensemble of 10,00010{,}000 independent trajectories of the L96 system:
  • The system dimension is set to $N=40$.
  • Initial conditions are generated by perturbing the equilibrium state $x_{i}=F$ with small Gaussian noise:
    $\mathbf{x}_{0}=F\cdot\mathbf{1}+\epsilon\cdot\mathcal{N}(0,I),\quad\epsilon=10^{-3}$
  • Each trajectory is simulated for a total time of 15 seconds.
  • To remove transient behavior and capture steady-state dynamics, the first 10 seconds (burn-in period) are discarded.
  • The remaining 5 seconds of simulation are retained, corresponding to $500+1=501$ time steps with $\Delta t=0.01$.
The resulting dataset has shape (10,000,501,40)(10{,}000,501,40), where each sample corresponds to a trajectory of 40-dimensional state vectors over 501 time steps. This dataset captures the spatiotemporal evolution of the system from a diverse set of initial conditions and serves as a challenging testbed for learning nonlinear operator mappings.
@@
Figure 14: Complete spatio-temporal field reconstruction comparison between DeepONet and FEDONet.
Figure 14 displays two representative hold-out trajectories of the L96 system. For each sample we plot the initial state, the reference spatio-temporal evolution, predictions from the DeepONet and FEDONet, together with the corresponding absolute-error fields.
  • Phase propagation. Both networks reproduce the diagonal stripe pattern that characterises L96 wave propagation, preserving phase speed and wavelength over the 500-step horizon. No visible drift or amplitude blow-up appears in either solution.
  • Error magnitude. The residual maps reveal only modest differences: in Sample 1 the RMS error of the vanilla model is $\mathcal{O}(10^{-1})$ while FEDONet trims it to the lower end of the same decade; Sample 2 shows the opposite ordering at isolated time–state patches. Overall, FEDONet yields a slightly lower global $\ell_{2}$ error but the gap is far narrower than that observed for the Poisson or Burgers benchmarks.
@@
  • Interpretation. Because the L96 solution manifold is dominated by a single, slowly varying spatio-temporal mode, a shallow fully connected trunk already provides a sufficiently rich basis. Injecting random Fourier features therefore offers limited extra benefit, consistent with the quantitative numbers in Table 1, where both models attain sub-25 % relative error.

### v2

@@ Abstract
[…]
@@ 1 Introduction
Partial differential equations underpin models ranging from turbulent‐flow dynamics and heat diffusion to electromagnetics and biological transport. Classical numerical solvers such as finite‐difference, finite‐element and spectral methods provide high‐fidelity solutions but incur a prohibitive computational burden when faced with high‐dimensional, multiscale or stiff systems [1, 2, 3]. Efforts to construct efficient surrogate models have a long history, spanning the universal approximation theorems for neural networks[4, 5, 6, 7], reduced‐order techniques based on proper‐orthogonal decomposition [8, 9], kernel Gaussian process surrogates [10], and mesh‐free radial‐basis schemes [11, 12]. However, extending these approaches to learn function‐to‐function mappings capable of generalizing across discretizations, boundary conditions, and input configurations has remained a long standing challenge.
The recent emergence of operator learning addresses this long-standing challenge by training neural networks that act directly on infinite-dimensional function spaces, enabling generalization across discretizations and grid resolutions. Deep Operator Networks [13] pioneered the branch–trunk architecture for learning nonlinear operators, while subsequent Neural Operator formulations [14, 15, 16] extended these ideas with rigorous operator–approximation guarantees and scalable architectures. Building on these foundations, Fourier Neural Operators [17] employ global spectral convolutions to capture nonlocal correlations efficiently, achieving remarkable speed-ups and accuracy gains in parametric PDE learning. Subsequent developments have further advanced this paradigm through multiwavelet representations [18, 19], convolutional kernels [20], multigrid tensorizations [21, 22], and geometry-aware spectral deformations [23], collectively enhancing scalability, resolution fidelity, and geometric adaptability.
Beyond architectural and kernel innovations, recent research has expanded operator learning along multiple frontiers. Efforts addressing solution irregularities have focused on handling discontinuities [24], incorporating derivative supervision for enhanced gradient fidelity [25], and enabling learning on irregular meshes through graph-based trunk networks [26]. Adaptive and multi-resolution strategies have been introduced to improve sampling efficiency [27], while hybrid encoder-decoder frameworks leveraging U-Net and wavelet representations further enhance multi-scale expressivity [28, 29]. To embed physical constraints and quantify uncertainty, the paradigm of Physics-Informed Neural Networks (PINNs) [30, 31] has been generalized into Physics-Informed Neural Operators [32, 33] and related variational or pseudo-physics extensions [34, 35]. Parallel efforts have explored latent and invertible operator formulations to achieve data-efficient and bidirectional inference [36, 37, 38, 39]. Transformer based attention mechanisms have also been adapted for operator learning [40, 41, 42, 43, 44] and for robust geometry generalization [45]. Together, these advances have enabled high-fidelity, climate-scale surrogates such as FourCastNet [46] and its amortized Fourier Neural Operator successor [47], underscoring the maturity and scalability of the operator-learning paradigm.
Despite the recent operator-learning architectures, the trunk network of conventional DeepONets remains a shallow multilayer perceptron. This design often struggles to represent the oscillatory and multi-scale structures that characterize nonlinear PDE solutions, resulting in reduced sample efficiency and reconstruction accuracy. While recent alternatives such as graph-based [26] or wavelet-enhanced U-Net trunks [28] offer improved expressivity, they introduce significant architectural complexity and domain-specific preprocessing requirements. In contrast, random Fourier features have proven remarkably effective at encoding high-frequency information in standard deep networks [48], yet their potential within operator-learning frameworks remains largely unexplored, where the trunk is typically implemented as a simple fully connected MLP. To address this gap, we propose the Fourier-Embedded Deep Operator Network (FEDONet), which augments the DeepONet trunk by replacing raw spatial coordinates with a compact random Fourier-feature embedding [48]. This embedding implicitly introduces a spectrum of sinusoidal basis functions, enriching the trunk’s representational capacity with negligible additional parameters. FEDONet retains the original branch–trunk separation, integrates seamlessly with existing DeepONet implementations, and incurs no runtime overhead beyond a single matrix multiplication.
We formally introduce the Fourier embedding of trunk coordinates and analyze its theoretical implications through the lens of the operator neural tangent kernel, demonstrating that the embedding strictly enlarges the approximation class of standard DeepONets. A comprehensive benchmark is carried spanning seven canonical PDE families - steady Poisson, viscous Burgers, Allen–Cahn, Kuramoto–Sivashinsky, Eikonal, Lorenz–63, and Lorenz–96. By embedding a principled spectral bias into the trunk network, FEDONet bridges the performance gap to frequency-domain operators, while preserving the architectural simplicity, flexibility, and locality benefits of the original DeepONet formulation.
Section 2 presents the formulation of the Fourier-embedded Deep Operator Network (FEDONet), detailing the embedding construction, its theoretical connection to the operator neural tangent kernel, and its implications for approximation capacity. Section 3 describes the benchmark setup and implementation details across the seven PDE families considered. Section 4 reports quantitative and spectral analyses of FEDONet performance, including convergence behavior, and comparisons with baseline operator models. Finally, Section 5 summarizes the main findings and outlines directions for future research in spectral operator learning and physics-constrained network design.
@@ 2.1 Fourier Embeddings as Spectral Preconditioners
Standard MLPs exhibit a well-known spectral bias - they learn low-frequency components faster than high-frequency ones [49]. This bottleneck is particularly harmful in operator learning problems involving sharp gradients, multiscale dynamics, or oscillatory fields.
To address this, we propose injecting fixed Fourier embeddings into the trunk network input illustrated in Fig. 1. Given a coordinate $\zeta\in\mathbb{R}^{d}$, we define a randomized Fourier feature map:
@@
where $B\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ - enriching its spectral content.
The FEDONet prediction then becomes:
@@
effectively lifting $\zeta$ into a high-dimensional basis of sinusoidal functions. By Bochner’s theorem [50], such embeddings approximate shift-invariant kernels - implicitly kernelizing the trunk network. This expands its representational capacity, enabling better learning of high-frequency targets.
@@ 2.2 Training Objective
We adopt a supervised learning framework to train the proposed FEDONet model. The objective is to approximate a target operator $G$ by minimizing the prediction error over a dataset of paired input-output functions. Each dataset element consists of a function $u^{(i)}\in\mathcal{U}$ and corresponding evaluations of the output $s^{(i)}_{j}=G(u^{(i)})(\zeta^{(i)}_{j})$ at query locations $\zeta^{(i)}_{j}\in\mathcal{Y}$.
@@
Model parameters are optimized using mini-batch stochastic gradient descent with the Adam optimizer. Learning rate schedules are selected based on the problem characteristics to ensure stable and efficient convergence. Each mini-batch consists of randomly sampled function-query pairs $(u,\zeta)$, enabling scalable and efficient training across diverse domains.
@@ 3 Results
We constructed a synthetic dataset comprising numerical solutions to the canonical two-dimensional Poisson equation. The governing equation is
@@
subject to homogeneous Dirichlet boundary conditions,
$u(x,y)\big|_{\partial\Omega}=0,$ (10)
where $u(x,y)$ denotes the potential field and $f(x,y)$ represents a prescribed source distribution.
A total of N=10, [^0] N=10{,}000 distinct source fields were synthesized using Gaussian random fields (GRFs) with spectral parameters $\alpha=3$ and $\tau=3$,
$f\sim\text{GRF}(\alpha=3,\tau=3),$ (11)
@@
on a uniform $128\times 128$ grid. For each realization of $f$, the corresponding solution $u$ was obtained by numerically solving the Poisson equation using a standard five-point finite-difference discretization of the Laplacian operator. The spatial domain $\Omega=[0,1]^{2}$ was thus transformed into a structured algebraic system,
$A\mathbf{u}=\mathbf{b},$ (12)
where $\mathbf{u},\mathbf{b}\in\mathbb{R}^{16384}$ denote the vectorized forms of the solution and source fields, respectively. The coefficient matrix $A$ was defined using a conventional five-point stencil with Dirichlet conditions enforced by setting boundary rows to the identity and zeroing the corresponding entries in $\mathbf{b}$. This process yielded 10,000 pairs of input-output fields suitable for training, validation, and testing of neural operator architectures.
Figure 2: Comparison of predicted and reference solution fields for the two-dimensional Poisson equation.
Figure 2 shows a representative test case from the Poisson dataset. The top row depicts the input forcing, reference solution, prediction from the baseline DeepONet, and the corresponding pointwise error, while the bottom row presents results from the proposed Fourier-Embedded DeepONet (FEDONet). FEDONet reproduces the reference field with sharper gradients and more accurate spatial localization of extrema, whereas the baseline model exhibits noticeable blurring and displacement of high-magnitude regions. The residual field of the vanilla DeepONet fluctuates within $\pm 1.5\times 10^{-3}$ and displays fine-scale artifacts, while FEDONet reduces the range to $\pm 3\times 10^{-4}$ and yields a smoother, spatially coherent error distribution, corresponding to an approximate fivefold improvement in reconstruction accuracy.
Spectral accuracy was assessed using the angle-integrated energy spectrum $E_{k}$ (Fig. 3). Both models reproduce the reference spectrum at large scales ($k\lesssim 5$), but differences arise in the intermediate and dissipative ranges. The baseline DeepONet consistently overestimates energy for $k\gtrsim 8$, resulting in a spurious buildup of high-wavenumber content that mirrors the oscillatory residuals seen in physical space. In contrast, FEDONet closely follows the reference spectrum across the entire range, deviating by less than a factor of two near the Nyquist limit and maintaining a physically consistent energy decay.
@@
Taken together, the spatial and spectral analyses demonstrate that the Fourier embedding enhances the trunk network’s ability to capture high-frequency features and improves overall stability. FEDONet achieves lower pointwise errors while eliminating the artificial spectral plateau that characterizes conventional fully connected trunks.
@@ 3.2 Burgers’ Equation
Figure 5: Best-performing test sample for Burgers’ Equation: Ground truth vs. DeepONet vs. FEDONet prediction.
@@
Figure 6: Temporal Evolution of Pointwise Error and Energy Spectrum for Burgers’ Equation
@@
Figure 7: Training loss convergence for the 1D Burgers’ Equation - Comparison of Vanilla and Fourier Embedded DeepONets trained with varying batch sizes.
@@ 3.3 Lorenz–63 System
A dataset of 10,00010{,}000 trajectories is generated using the classic Runge-Kutta method (RK45), each evaluated at 1000 uniformly spaced time steps. The input to the DeepONet branch network is the scalar $x_{0}$, and the trunk network is queried at time coordinates $t\in[0,3]$. The goal is to accurately recover the full state trajectory, including lobe transitions and fine-scale oscillations characteristic of chaotic dynamics.
@@
Finally, Figure 12 compares the spectral content of the predicted and reference trajectories. The vanilla DeepONet fails to capture high-frequency modes, resulting in a smoothed trajectory that diverges from the correct dynamics. FEDONet, equipped with randomized Fourier features, preserves the broadband energy spectrum more accurately - an essential factor in recovering the fine-scale structure and temporal richness of chaotic systems.
@@ 3.4 Eikonal Equation
Figure 13: Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry for Eikonal Equation. Relative $L^{2}$ error: $2.728\times 10^{-2}$ for DeepONet, $0.912\times 10^{-2}$ for FEDONet.
@@ 3.5 Lorenz-96 Chaotic System
To assess model performance on nonlinear dynamical systems, we consider the Lorenz-96 (L96) model, a prototypical chaotic system widely used for testing data-driven surrogates in climate modeling and nonlinear forecasting. The governing equations are
$\frac{dx_{i}}{dt}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,\quad i=1,2,\dots,N,$ (26)
@@
subject to periodic boundary conditions $x_{i-N}=x_{i}=x_{i+N}$. Here, $x_{i}$ denotes the state of the $i$-th variable, $N$ is the system dimension, and $F$ is a constant forcing parameter that determines the level of nonlinearity and chaos. For $F>8$, the system exhibits strong chaotic behavior; in this study, we set $F=4.0$ to obtain moderately nonlinear yet stable dynamics.
@@
The system is integrated using a classical fourth-order Runge–Kutta (RK4) scheme with a fixed time step $\Delta t=0.01$:
$\mathbf{x}_{n+1}=\mathbf{x}_{n}+\frac{\Delta t}{6}(k_{1}+2k_{2}+2k_{3}+k_{4}),$ (27)
where $k_{1},k_{2},k_{3},k_{4}$ are the intermediate RK stages computed from the Lorenz–96 derivatives.
A dataset of 10,00010{,}000 independent trajectories is generated for training and evaluation. Each trajectory is initialized as
$\mathbf{x}_{0}=F\cdot\mathbf{1}+\epsilon\,\mathcal{N}(0,I),\quad\epsilon=10^{-3},$
with $N=40$ and integrated for a total physical time of 15 s. The initial 10 s are discarded as a transient, and the remaining 5 s (501 time steps) are retained to capture quasi-stationary behavior. The resulting dataset has dimensions (10,000,501,40)(10{,}000,501,40), representing 10,000 spatio-temporal trajectories of 40-dimensional state vectors.
@@
Figure 14: Lorenz-96 - Complete spatio-temporal field reconstruction comparison between DeepONet and FEDONet.
Figure 14 shows representative test trajectories from the L96 dataset. Each panel depicts the initial condition, reference evolution, predictions from the baseline DeepONet and the proposed Fourier-Embedded DeepONet (FEDONet), and their respective absolute-error fields. Both models reproduce the characteristic diagonal stripe patterns associated with wave-like phase propagation, maintaining amplitude and wavelength stability over the full 500-step forecast horizon. No visible drift or blow-up is observed in either solution.
Quantitatively, the residual fields indicate comparable performance. In the first trajectory, the vanilla DeepONet attains a root-mean-square (RMS) error on the order of $\mathcal{O}(10^{-1})$, while FEDONet slightly reduces this magnitude; in the second trajectory, the ordering reverses locally, but the overall difference remains small. On average, FEDONet achieves marginally lower global $\ell_{2}$ errors, though the improvement is less pronounced than in the Poisson or Burgers benchmarks.
@@
This outcome reflects the inherent structure of the L96 dynamics: the solution manifold is dominated by a slowly varying, low-dimensional spatio-temporal mode, which a shallow fully connected trunk can already approximate effectively. Consequently, introducing random Fourier features provides limited additional benefit, consistent with the quantitative results in Table 1, where both models achieve sub-25% relative error.